OpenStudy (anonymous):

Solve Initial Value Problem? dy/dx=y/x and y=2 when x=2

OpenStudy (usukidoll):

oooo differential equations? separable equations?

OpenStudy (usukidoll):

ok so we need x's on one side and y's on the other.. in the form of h(y) dy = f(x) dx

OpenStudy (usukidoll):

so multiply the y/x by dx and divide the y... this is so hard without draw!

OpenStudy (anonymous):

Okay done

OpenStudy (usukidoll):

http://www.twiddla.com/1474663

OpenStudy (usukidoll):

we need a y not a ln y so I think we need to e to the everything

OpenStudy (anonymous):

my bad i didn't even see you typing.

OpenStudy (usukidoll):

WHAT THE HECK?

OpenStudy (anonymous):

lol :P

OpenStudy (usukidoll):

I thought antiderivative of 1/y 1/x is ln y and ln x? :P

OpenStudy (usukidoll):

ufkodfjdskfdslj;sdaf ljdfslkj I have to finish my latex or I'm screwed...

OpenStudy (anonymous):

i did dy/dx = x/y good thing you're here :P

OpenStudy (usukidoll):

yeah but we need the equation in the form of h(y) dy = f(x) dx

OpenStudy (anonymous):

Well now I'm confused

zepdrix (zepdrix):

:o

OpenStudy (usukidoll):

so ... dx needs to go to the right the y oh wait. ugh that could be me screwing it up by accident... it would be let me redraw this.

zepdrix (zepdrix):

Equation and draw tool are still broken :c le sigh...

OpenStudy (usukidoll):

yeah I effed up by accident sorry long day at school tired D:

OpenStudy (usukidoll):

http://www.twiddla.com/1474663

zepdrix (zepdrix):

Ooo neat tool :O Never seen this before!

OpenStudy (anonymous):

Why x/y ?

OpenStudy (usukidoll):

there whew..........

OpenStudy (usukidoll):

this is a separable equation... please go to the link for the drawing of your problem

OpenStudy (usukidoll):

you need all y's to the left and all x's to the right

zepdrix (zepdrix):

The original problem was y/x silly D:

OpenStudy (usukidoll):

OMG just see my drawing D:<

zepdrix (zepdrix):

But..... :(

zepdrix (zepdrix):

You started with dy/dx = x/y the original problem is dy/dx = y/x .... :C

OpenStudy (usukidoll):

ughhhhh sorry my bad...oh man second time fine third time is the charm as they say

zepdrix (zepdrix):

XD

zepdrix (zepdrix):

Why is the cutting of the bottom quarter of the page on twiddla? Am I the only one having that problem? :[ Hmm weird..

OpenStudy (usukidoll):

****** I was right earlier with the lns who is trying to screw me up ? D:

OpenStudy (anonymous):

Sigh...

zepdrix (zepdrix):

Yes you were on the right track the first time it appears :o

OpenStudy (usukidoll):

yeah someone threw me off. >:/

OpenStudy (usukidoll):

@Euler271 DID IT >:O

zepdrix (zepdrix):

And now I'm stealing your thunder XD Sneaky green text lol

OpenStudy (anonymous):

change to (1/y) dy/dx = (1/x) Then integrate both sides with respect to x.

zepdrix (zepdrix):

This twiddla is way cool. Imma have to use this til we get our site utilities up and working.

OpenStudy (anonymous):

So zepdrix your answer is correct?

zepdrix (zepdrix):

I think so :o It got erased t.t pony got mad

OpenStudy (anonymous):

I wrote it all down

zepdrix (zepdrix):

Did you understand the exponentiation step she was doing? :o That part can be a little tricky.

zepdrix (zepdrix):

Particularly, did you understand why the C is now being multiplied instead of added?

OpenStudy (usukidoll):

exponential rules

OpenStudy (anonymous):

Yeah, I got it:) Thanks for the help. Phew. I need some sleep :P